### v1

# REST: Diffusion-based Real-time End-to-end Streaming Talking Head Generation via ID-Context Caching and Asynchronous Streaming Distillation

Abstract
Diffusion models have significantly advanced the field of talking head generation. However, the slow inference speeds and non-autoregressive paradigms severely constrain the application of diffusion-based THG models. In this study, we propose REST, the first diffusion-based, real-time, end-to-end streaming audio-driven talking head generation framework. To support real-time end-to-end generation, a compact video latent space is first learned through high spatiotemporal VAE compression. Additionally, to enable autoregressive streaming within the compact video latent space, we introduce an ID-Context Cache mechanism, which integrates ID-Sink and Context-Cache principles to key-value caching for maintaining temporal consistency and identity coherence during long-time streaming generation. Furthermore, an Asynchronous Streaming Distillation (ASD) training strategy is proposed to mitigate error accumulation in autoregressive generation and enhance temporal consistency, which leverages a non-streaming teacher with an asynchronous noise schedule to supervise the training of the streaming student model. REST bridges the gap between autoregressive and diffusion-based approaches, demonstrating substantial value for applications requiring real-time talking head generation. Experimental results demonstrate that REST outperforms state-of-the-art methods in both generation speed and overall performance.
Machine Learning, ICML
1 Introduction
Audio-driven Talking Head Generation (THG) aims to generate realistic talking head videos conditioned on input speech and reference images (Chen et al., 2020), demonstrating significant research and application value in domains such as remote education (Gowda et al., 2023), virtual reality (Zhen et al., 2023), and film production (Song et al., 2023). The lip-sync accuracy and naturalness of the generated videos serve as fundamental benchmarks for model performance (Song et al., 2023). Furthermore, due to growing demands of real-time applications like human-computer interaction, the inference speed of THG models is now garnering increasing research attention (Zhen et al., 2023).
Recent advances in talking head generation have been significantly driven by the introduction of diffusion models (Croitoru et al., 2023). Compared to earlier approaches (Prajwal et al., 2020), diffusion-based THG methods (Wang et al., 2025b, c) demonstrate substantial improvements in both lip-sync accuracy and visual naturalness. Despite the improvements, two major limitations remain in current diffusion-based THG systems. First, end-to-end diffusion models suffer from slow inference speeds, typically requiring tens to hundreds of seconds to generate a five-second video (Meng et al., 2025). Second, most existing diffusion-based models operate in a non-streaming manner, which further increases latency for end users (Wang et al., 2025b). Conversely, autoregressive (AR) THG methods (Chu et al., 2025) typically generate intermediate facial motion representations from speech in a sequential manner, which are then rendered into video via a separate rendering module, achieving lower latency and supporting streaming inference beyond diffusion-based methods. However, the output video quality of two-stage AR methods is often limited by the expressiveness of the generated motion representations, resulting in generally inferior performance compared to end-to-end diffusion-based approaches.
To bridge the gap between diffusion-based and AR-based THG methods in addressing key challenges in the field of THG, in this research we introduce REST, the first diffusion-based real-time end-to-end streaming audio-driven talking head generation framework. We first utilize a Variational Autoencoder (VAE) with high spatiotemporal compression to derive a compact latent video representation, reducing the computational burden of end-to-end diffusion. Inspired by AR methods, we integrate a novel ID-Context Cache mechanism into our diffusion transformer backbone. By combining ID-Sink and Context-Cache, ID-Context Cache enables chunk-by-chunk AR inference while enhancing identity fidelity and inter-frame temporal consistency. Furthermore, to enhance temporal smoothness and generation quality in streaming generation, an Asynchronous Streaming Distillation (ASD) scheme is proposed, which leverages latent spatiotemporal context generated by a non-streaming teacher model to supervise the streaming student model.
Figure 1: Performance benchmarking of the SOTA THG models.
In summary, our contributions are as follows:
  • We propose REST, the first diffusion-based real-time end-to-end streaming talking head generation model that bridges the gap between diffusion-based and AR-based THG methods.
  • We introduce a novel ID-Context Cache mechanism that incorporates ID-Sink and Context KV Cache principles, enabling AR-style inference while maintaining identity fidelity and inter-frame temporal consistency.
  • We further propose an Asynchronous Streaming Distillation scheme for streaming teacher-student supervised learning, leveraging non-streaming spatiotemporal latent context to supervise the streaming student model, achieving better streaming performance.
2 Related Work
2.1 Audio-driven Talking Head Generation
Audio-driven Talking Head Generation (THG) is a cross-modal audio-to-video synthesis task aimed at generating realistic talking head videos conditioned solely on a reference image and speech audio input. Early efforts in Audio-driven THG (Prajwal et al., 2020; Stypułkowski et al., 2024) primarily focused on achieving lip-sync between the generated video and the input speech. Subsequent work (Wang et al., 2021; Zhang et al., 2023) expanded beyond lip-sync to improve the overall naturalness of the output, regrading natural facial expressions and motions. Recently, the success of diffusion models (Nichol & Dhariwal, 2021; Song et al., 2020) in image and video generation (Croitoru et al., 2023) has revolutionized THG. End-to-end THG frameworks built upon pretrained image or video diffusion models have demonstrated superior performance on lip-sync and motion naturalness. However, these improvements come at the cost of substantially increased inference latency. Moreover, to ensure better temporal consistency, most current diffusion-based THG models typically adopt non-autoregressive paradigms that process long video chunks. As a result, users must wait for the entire chunk to be synthesized before viewing, which hinders real-world deployment. To address this limitation, we propose an end-to-end streaming THG framework based on ID-Context Cache, along with a novel streaming teacher-student learning pipeline that achieves identity fidelity and temporal consistency comparable to non-streaming models.
2.2 Fast Talking Head Generation
Talking Head Generation differs from general video generation in its stronger demand for real-time performance, particularly in applications such as human-computer interaction and game production (Zhen et al., 2023). Conventional efficient THG often adopts a two-stage generation paradigm. In this approach, low-dimensional 2D or 3D motion representations (Li et al., 2025; Liu et al., 2024), are utilized as video driving signals, and a renderer is employed to generate videos from these representations, accelerating inference by converting the high-dimensional latent generation task into a low-dimensional motion generation problem. Motion representation generation is primarily approached via diffusion or AR schemes. Motion-space diffusion methods such as AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025) utilize lightweight diffusion models to predict motion representations, while AR-based approaches (Chu et al., 2025) generate motions in a context-aware manner to enable streaming capabilities. However, two-stage methods are bottlenecked by the fidelity of generated motion representations, generally yielding naturalness inferior to end-to-end approaches (Wang et al., 2025a; Ji et al., 2025). READ (Wang et al., 2025a) represents the first attempt at real-time end-to-end talking head generation, achieving audio-visual alignment in a highly compressed latent space, but its non-autoregressive design still fails to support streaming real-time synthesis. In this work, we propose the first diffusion-based end-to-end streaming THG model that successfully bridges AR and diffusion paradigms.
Figure 2: The overall framework of REST. During training, the non-streaming teacher model is first pre-trained by asynchronous chunk-wise scheduler to provide a solid performance upper bound, shown in ①. Then the streaming student model with ID-Context Cache is trained under the guidance of the streaming teacher by ASD, shown in ②. The principle of ASD is demonstrated in ③
3 Methods
The total framework of REST is shown in Fig. 2. Sec. 3.1 outlines the necessary preliminaries relevant to our work. Sec. 3.2 details the proposed model architecture with ID-Context Cache. Sec. 3.3 focuses on the model training methodology guided by ASD. And the final section gives an introduction on the AR-style inference methodology of the parametrized streaming model.
3.1 Preliminaries
Task Definition. Audio-driven THG task can be formulated as a mapping from a reference image $\mathit{\boldsymbol{I}_{\text{ref}}}\in\mathbb{R}^{H\times W\times 3}$ and speech $\mathit{\boldsymbol{S}_{1:F_{a}}}$ to an output video $\mathit{\boldsymbol{X}}\in\mathbb{R}^{H\times W\times 3\times F_{v}}$, where $F_{a}$ and $F_{v}$ denotes the length of speech and output video.
Flow Matching. The central idea of Flow Matching (FM) (Lipman et al., 2022) is to learn a continuous-time vector field ${\mathit{\boldsymbol{v}}}({\mathit{\boldsymbol{Z}}}(t),t)$ that transports samples from a simple noise distribution ${\mathit{\boldsymbol{Z}}}(t)$ to the target data distribution ${\mathit{\boldsymbol{Z}}}(0)$ (Lipman et al., 2024; Dao et al., 2023):
${{\mathit{{d}}}{\mathit{\boldsymbol{Z}}}(t)}={\mathit{\boldsymbol{v}}}({\mathit{\boldsymbol{Z}}}(t),t){dt}$ (1)
The forward process of FM defines a probability path from the original distribution ${\mathit{\boldsymbol{Z}}}(0)$ to ${\mathit{\boldsymbol{Z}}}(t)$. The process can be formulated when using Gaussian probability paths to add synchronous Gaussian noise at timestep $t$ to ${\mathit{\boldsymbol{Z}}}(0)$:
${\mathit{\boldsymbol{Z}}}(t)=(1-t){\mathit{\boldsymbol{Z}}}(0)+t\boldsymbol{\epsilon},~\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I})$ (2)
The training objective of FM is for the model $\theta$ to learn the correct vector field ${\mathit{\boldsymbol{u}}}({\mathit{\boldsymbol{Z}}}(t),t)$, as follows:
[…]
where ${\mathit{\boldsymbol{u}}}({\mathit{\boldsymbol{Z}}}(t),t)$ is the target gound-truth vector field.
3.2 Streaming Diffusion with ID-Context Cache
To achieve real-time streaming generation, we propose a framework comprising a temporal VAE and a DiT backbone integrated with the proposed ID-Context Cache mechanism.
Temporal VAE for compact latent space. For real-time efficiency in end-to-end diffusion, a compact latent space of videos is learned via a temporal VAE, which achieves a high compression ratio of 32×32×8 pixels per token, as inspired by LTX-Video (HaCohen et al., 2024). The principle can be formulated as follows:
${\mathit{\boldsymbol{Z}}}(0)=\mathcal{E}_{\text{V}}({\mathit{\boldsymbol{X}}}(0)),~{\hat{\mathit{\boldsymbol{X}}}}(0)=\mathcal{D}_{\text{V}}({\mathit{\boldsymbol{Z}}}(0))$ (4)
where ${\mathit{\boldsymbol{X}}}(0)\in\mathbb{R}^{H\times W\times F\times D_{\text{v}}}$ denotes the video sequence, and ${\mathit{\boldsymbol{Z}}}(0)\in\mathbb{R}^{h\times w\times f\times d_{\text{v}}}$ are the compressed video latents. $\mathcal{E}_{\text{V}}$ and $\mathcal{D}_{\text{V}}$ denotes the encoder and decoder of VAE.
SpeechAE for compact speech latent space. To facilitate cross-modal alignment within the compact latent space, we employ the SpeechAE architecture (Wang et al., 2025a) to perform synchronous temporal compression on the input speech embedding $\mathit{\boldsymbol{S}_{1:F}}$. SpeechAE first integrates a Whisper-tiny encoder (Radford et al., 2023) for speech embedding extraction, as described below:
$\mathit{\boldsymbol{S}_{1:F}}=\mathcal{E}_{\text{Whisper}}(\mathit{\boldsymbol{A}_{1:F_{a}}})$ (5)
Then the corresponding compact speech latent embedding $\mathit{\boldsymbol{E}\in\mathbb{R}^{f\times h_{w}\times d_{\text{A}}}}$ are generated by the SpeechAE encoder from $\mathit{\boldsymbol{S}}$, and reconstructed to $\hat{\mathit{\boldsymbol{S}}}$ through the decoder:
${\mathit{\boldsymbol{E}}}=\mathcal{E}_{\text{A}}({\mathit{\boldsymbol{S}}}),~\hat{{\mathit{\boldsymbol{S}}}}=\mathcal{D}_{\text{A}}({\mathit{\boldsymbol{E}}})$ (6)
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims.
ID-Context Cache for autoregressive diffusion. To enable autoregressive, context-aware generation in audio-to-video diffusion, we introduce a novel DiT-based backbone that implements the proposed ID-Context Cache mechanism. Our Streaming A2V-DiT framework comprises 28 transformer blocks, each integrating three core modules: Self-Attention with ID-Context Cache, 3D Full-Attention for text conditioning, and Frame-level 2D Cross-Attention for audio conditioning. Theoretically, the ID-Context Cache reconfigures the attention topology to enable chunk-by-chunk autoregressive inference, simultaneously enforcing ID fidelity and temporal causality. This is realized through the two following theoretically motivated components.
ID-Sink Principle: ID-Sink models the Key-Value (KV) embeddings of the reference image as persistent sinks across all chunks, establishing a global semantic anchor to preserve identity fidelity. Context-Cache Principle: Context Cache approximates a continuous temporal flow by extending the effective temporal receptive field. It reconstructs local temporal dependencies by concatenating the KV embeddings of the preceding chunk with the current KV embedding during self-attention, thereby mitigating boundary discontinuities.
The mathematical formulation of the ID-Context Cache is defined as follows. For the current chunk, the Query, Key, and Value at each layer are computed as formulated below:
$\mathit{\boldsymbol{Q}_{ij}}=\mathit{\boldsymbol{W}_{j}^{Q}}\cdot\mathit{\boldsymbol{H}_{ij}}$ (7)
$\mathit{\boldsymbol{K}_{ij}^{c}}=\mathit{\boldsymbol{W}_{j}^{K}}\cdot\mathit{\boldsymbol{H}_{ij}}$ (8)
$\mathit{\boldsymbol{V}_{ij}^{c}}=\mathit{\boldsymbol{W}_{j}^{V}}\cdot\mathit{\boldsymbol{H}_{ij}}$ (9)
where $i$ denotes the $i$-th chunk and $j$ denotes the $j$-th block. Then, the causal dependency chain is reconstructed by concatenating the static ID anchor with the dynamically updated contextual cache from the preceding chunk, following the proposed ID-Sink and Context Cache principles:
$\mathit{\boldsymbol{K}_{ij}}=[\mathit{\boldsymbol{K}}_{0j}^{c}\mathbin{\|}\mathit{\boldsymbol{K}}_{(i-1)j}^{c}\mathbin{\|}\mathit{\boldsymbol{K}}_{ij}^{c}]$ (10)
$\mathit{\boldsymbol{V}_{ij}}=[\mathit{\boldsymbol{V}}_{0j}^{c}\mathbin{\|}\mathit{\boldsymbol{V}}_{(i-1)j}^{c}\mathbin{\|}\mathit{\boldsymbol{V}}_{ij}^{c}]$ (11)
The self-attention is then operated based on the refined causal dependency chain, as follows:
$\mathit{\boldsymbol{H}}_{i(j+1)}=\mathit{\boldsymbol{H}}_{ij}+\text{Softmax}\left(\frac{\mathit{\boldsymbol{Q}}_{ij}\cdot\mathit{\boldsymbol{K}}_{ij}^{\text{T}}}{\sqrt{d_{k}}}\right)\cdot\mathit{\boldsymbol{V}}_{ij}$ (12)
Regarding cross-modal audio conditioning, the speech latent code ${\boldsymbol{E}}=[\mathit{\boldsymbol{e}}_{0}\mathbin{\|}\mathit{\boldsymbol{e}}_{1}\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{e}}_{k}]$ temporally aligned with the video chunks. By enforcing a frame-level spatial cross-attention with a temporal receptive field of chunk length, we achieve precise streaming audio-visual alignment within the compact latent space, formulated as:
$\mathit{\boldsymbol{H}}_{ij}^{\text{A}}=\mathit{\boldsymbol{H}}_{ij}+\text{CrossAttn}\left(\mathit{\boldsymbol{H}}_{ij},{\mathit{\boldsymbol{e}}_{j}}\right)$ (13)
where $\mathit{\boldsymbol{H}}_{ij},\mathit{\boldsymbol{H}}_{ij}^{\text{A}}\in\mathbb{R}^{h\times w\times f\times d}$ denotes the unpatchified hidden states before and after frame-level spatial cross-attention of $i\!-\!\text{th}$ attention block. Our proposed design enables the AR-style streaming generation of video latent chunks that are synchronized with the corresponding speech conditions.
3.3 Training with Asynchronous Streaming Distillation
To mitigate autoregressive error propagation and enforce inter-chunk temporal consistency in chunk-by-chunk streaming diffusion, we introduce the Asynchronous Streaming Distillation (ASD) paradigm. Theoretically, ASD aims to distill global temporal priors from a non-streaming teacher into a streaming student through asynchronous noise conditioning.
Asynchronous Non-streaming Teacher Training. To emulate streaming dynamics within a non-streaming framework for establishing a theoretical upper bound of temporal consistency, we employ a chunk-wise asynchronous noise scheduler to guide the training of the non-streaming teacher. The teacher model receives the full latent sequence as input. We first concatenate the reference frame to the front of the initial video latents to provide speaker ID, as follows:
$\mathit{\boldsymbol{z}}_{\text{R}}=\mathcal{E}_{\text{V}}(\mathit{\boldsymbol{I}_{\text{ref}}}),{\mathit{\boldsymbol{Z}}}(0)=[\mathit{\boldsymbol{z}}_{\text{R}}\mathbin{\|}\mathit{\boldsymbol{z}}_{1}(0)\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{z}}_{k}(0)]$ (14)
Then the asynchronous add-noise process operates by applying a chunk-wise noise instance to each chunk, as follows:
$\boldsymbol{t}\!=\![0\mathbin{\|}\underbrace{t_{1},\cdots,t_{1}}_{\text{1-th Chunk}}\mathbin{\|}\underbrace{t_{2},\cdots,t_{2}}_{\text{2-th Chunk}}\mathbin{\|}\cdots\mathbin{\|}\underbrace{t_{k},\cdots,t_{k}}_{\text{k-th Chunk}}]$ (15)
$\!{\mathit{\boldsymbol{Z}}}(\boldsymbol{t})\!=\!(\boldsymbol{1}-\boldsymbol{t})\odot{\mathit{\boldsymbol{Z}}}(0)+\boldsymbol{t}\odot\boldsymbol{\epsilon}$ (16)
The final training objective of the teacher network parameters $\mathcal{T_{\theta}}$ is as follows, conditioned on the audio latents $\boldsymbol{E}$:
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-Z(0)$ (17)
[…]
where $\boldsymbol{v}$ denotes the correct vector field under the Gaussian probability path of the asynchronous add-noise process.
Streaming Student Training. Subsequently, to emulate inference-time causal dynamics, the student model is trained via a sequential chunk-wise paradigm. We apply the same add-noise paradigm as teacher model to obtain noisy latents ${\mathit{\boldsymbol{Z}}}(t)$ from the initial latents ${\mathit{\boldsymbol{Z}}}(0)$, which are then partitioned into chunks $\widetilde{\mathit{\boldsymbol{Z}}}(t)=[\mathit{\boldsymbol{z}}_{\text{R}}\mathbin{\|}\mathit{\boldsymbol{z}}_{1}(t)\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{z}}_{k}(t)]$. These chunks are processed sequentially by the student model with the proposed ID-Context Cache mechanism to execute autoregressive generation. Also, the final training objective of the student network parameters $\mathcal{S_{\theta}}$ is as follows, conditioned on the audio latent chunks $\widetilde{\boldsymbol{E}}=[\mathit{\boldsymbol{e}}_{0}\mathbin{\|}\mathit{\boldsymbol{e}}_{1}\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{e}}_{k}]$:
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-Z(0)$ (19)
[…]
Teacher-Student Learning with ASD. The non-streaming teacher model leverages global attention to achieve superior temporal consistency, while the streaming student is limited by local causal attention under streaming constraints. Simply enlarging the Context Cache to broaden the receptive field leads to unsustainable computational costs. To address this fundamental limitation, we propose distribution matching between the non-streaming teacher and the streaming student, introducing global consistency constraints while adhering to streaming constraints. Specifically, we introduce a contrastive distribution loss that minimizes the distribution distance between teacher output flow $\hat{\mathit{\boldsymbol{v}}}=[\hat{\mathit{{v}}}_{0},\cdots,\hat{\mathit{{v}}}_{f}]$ and student output flow $\hat{\widetilde{\mathit{\boldsymbol{v}}}}=[\hat{\widetilde{\mathit{{v}}}}_{0},\cdots,\hat{\widetilde{\mathit{{v}}}}_{f}]$ at corresponding spatiotemporal positions, while maximizing the distance at different positions, as follows:
$\mathcal{L}_{\text{CON}}\!=\!-\frac{1}{f}\sum_{i=1}^{f}\mathrm{log}\!\left(\frac{\mathrm{exp}\left(\frac{\mathrm{sim}(\hat{\widetilde{\mathit{\boldsymbol{v}}}}_{i},\hat{\mathit{\boldsymbol{v}}}_{i})}{\tau}\right)}{\sum_{j=1,j\neq i}^{f}\mathrm{exp}\left(\frac{\mathrm{sim}(\hat{\widetilde{\mathit{\boldsymbol{v}}}}_{i},{\hat{\mathit{\boldsymbol{v}}}}_{j})}{\tau}\right)}\right)$ (21)
Furthermore, to enforce inter-chunk temporal consistency, we introduce a smoothness matching objective that minimizes the divergence in second-order optical flow between the outputs of teacher and student models at corresponding temporal positions. This regularization serves as a global consistency constraint, formulated as:
$\Delta v_{i}$ $=v_{i+1}-2v_{i}+v_{i-1}$ (22)
$\mathcal{L}_{\text{SMO}}$ $=\sum\nolimits_{i=1}^{f-1}\Delta v_{i}$ (23)
The final diffusion objective for the streaming student is formulated as a reconstruction loss regularized by teacher-student flow contrast and smoothness matching constraints:
$\mathcal{L}_{\text{SD}}=\mathcal{L}_{\mathcal{S}_{\theta}}+\mathcal{L}_{\text{CON}}+\mathcal{L}_{\text{SMO}}$ (24)
By aligning asynchronous teacher-student learning with the proposed constraints, global contextual knowledge from the non-streaming teacher can be effectively distilled into the streaming student, thereby enhancing the stability and temporal consistency in streaming generation.
3.4 AR-style Streaming Inference
Leveraging the ID-Context Cache and ASD strategies, the streaming model acquires the capability to condition current chunk synthesis on historical generative distributions. Based on this, we formulate an AR-style inference pipeline. The inference pipeline employs a dual-loop inference scheme, where the inner loop performs intra-chunk temporal diffusion and the outer loop executes inter-chunk recursion, progressively propagating historical information and ID anchors through ID-Context Cache. The resulting velocity flow is resolved to the latent space by an FM scheduler and reconstructed to pixel space via the temporal VAE decoder. During inference, we incorporate the Joint CFG (Wang et al., 2025a) mechanism for classifier-free guidance. The complete inference pipeline is detailed in Algorithm 1.
Algorithm 1 Asynchronous Reverse Process
Input :
Time schedule $\{T_{1},\cdots,T_{n}\}$ (where $T_{n}=0$);
Reference image $\boldsymbol{I}_{\text{ref}}$, $\boldsymbol{z}_{\text{R}}=\mathcal{E}_{\text{V}}(\boldsymbol{I}_{\text{ref}})$;
Speech latents $\boldsymbol{E}\in\mathbb{R}^{N\times h_{w}\times d_{a}}$;
Noise $\boldsymbol{\epsilon}=\{\boldsymbol{\epsilon}_{1},\!\cdots,\!\boldsymbol{\epsilon}_{N}\}\!\in\!\mathcal{N}(\mathbf{0},\mathbf{I})$, $\boldsymbol{Z}(T_{1})=\boldsymbol{\epsilon}$;
Zeros ID-Context Cache $\boldsymbol{C}=\{\boldsymbol{C}_{1},\cdots,\boldsymbol{C}_{n}\}$.
Output :
Generated latents $\boldsymbol{Z}(0)\in\mathbb{R}^{h\times w\times N\times d_{v}}$.
for $i\leftarrow 1$ to $k$ do
[…]
$\boldsymbol{Z}_{j}(T_{i+1}),\boldsymbol{C}_{j+1}\xleftarrow[\text{FM}]{\text{CFG}}\mathcal{S}_{\theta}\big(\boldsymbol{Z}_{j}(T_{i}),\boldsymbol{E}_{j},\boldsymbol{t}_{j},\boldsymbol{C}_{j}\big)$
   end for
end for
$\boldsymbol{Z}(0)\leftarrow\{\boldsymbol{Z}_{1}(T_{n}),\cdots,\boldsymbol{Z}_{k}(T_{n})\}$
// Updatereturn $\boldsymbol{X}(0):\boldsymbol{X}(0)=\mathcal{D}_{\text{V}}(\boldsymbol{Z}(0))$
4 Experiments and Results
4.1 Experimental Setup
Implementation Details. Experiments encompassing both training and inference are conducted on HDTF (Zhang et al., 2021) and MEAD (Wang et al., 2020) datasets. 95% data of both datasets is randomly allocated for training and the remaining 5% for testing, ensuring no overlap between the partitions. We employ a two-stage training strategy. In the first stage, the non-streaming teacher model is pre-trained at a resolution of $512\times 512$ pixels with a learning rate of $1\times 10^{-5}$. In the second stage, the streaming student model with ID-Context Cache is trained at a resolution of $512\times 512$ pixels and 121 frames, with a learning rate of $1\times 10^{-5}$ and a batch size of 1. All reported results use 8-step sampling and Joint-CFG with $\alpha=6.0$ unless specified. Both training and evaluation are performed on NVIDIA A100 GPUs.
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer, 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung & Zisserman, 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al., 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
Baselines. We benchmark our method against several SOTA open-source methods, including end-to-end diffusion methods such as Sonic (Ji et al., 2025), EchoMimic (Chen et al., 2025), Hallo (Xu et al., 2024), FantasyTalking (Wang et al., 2025c) and AniPortrait (Wei et al., 2024), as well as motion-space diffusion methods like AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025). All comparisons are conducted on the same device using identical test data with the same length of 4.84s (121 frames) to ensure fair evaluation.
4.2 Overall Comparisons.
As shown in Tab. 1, motion-space diffusion approaches such as Ditto and AniTalker generally demonstrate lower latency compared to end-to-end baselines. In contrast, our proposed end-to-end solution not only achieves a lower latency superior to all existing methods, but also pioneers streaming inference capabilities, which remains unattainable by existing alternatives. Beyond this efficiency, our approach achieves highly competitive performance, achieving either the best or second-best results across all metrics on both datasets, representing a significant breakthrough in real-time streaming talking head video generation.
Dataset Method Streaming Runtime(s) FID ($\downarrow$) FVD ($\downarrow$) E-FID ($\downarrow$) Sync-C ($\uparrow$) Sync-D ($\downarrow$)
HDTF Fantasy 896.089 16.489 315.291 1.232 5.138 10.349
Hallo 212.002 15.929 315.904 0.931 6.995 7.819
EchoMimic 124.105 18.384 557.809 0.927 5.852 9.052
Sonic 83.584 16.894 245.416 0.932 8.525 6.576
AniPortrait 76.778 17.603 503.622 2.323 3.555 10.830
Ditto 17.974 15.440 399.965 2.659 5.458 9.565
AniTalker 13.577 39.155 514.388 1.523 5.838 8.736
Ours 4.421 14.597 219.87 0.931 8.335 6.701
MEAD Fantasy 896.089 46.617 257.077 1.510 4.536 10.699
Hallo 212.002 52.300 292.983 1.171 6.014 8.822
EchoMimic 124.105 65.771 667.999 1.448 5.482 9.128
Sonic 83.854 47.070 218.308 1.434 7.501 7.831
AniPortrait 76.778 54.621 531.663 1.669 1.189 13.013
Ditto 17.974 45.403 349.860 1.941 5.199 9.595
AniTalker 13.577 95.131 621.528 1.553 6.638 8.184
Ours 4.421 46.540 237.521 1.064 7.632 7.573
Table 1: Overall comparisons on HDTF and MEAD. $\uparrow$” indicates better performance with higher values, while $\downarrow$” indicates better performance with lower values. The best results are bold, and the second-best results are underlined.
4.3 Ablation Studies
To assess the contribution of each component in our proposed method, several ablation studies are carried out.
Method FID ($\downarrow$) FVD ($\downarrow$) Sync-C ($\uparrow$)
Full ID‑Context 14.597 219.87 8.335
w/o ID-Sink 19.362 294.692 8.345
w/o Context-Cache 19.656 271.508 6.909
Table 2: Ablation results of ID-Context Cache on HDTF dataset.
Figure 3: Visualization of the ablation results of ID-Context Cache.
Ablation Studies on ID-Context Cache. To investigate the specific contributions of the ID-Sink and Context-Cache principles within our proposed ID-Context Cache methodology, we conduct an ablation study under different experimental configurations, as follows:
  • Full ID‑Context Cache: Our complete strategy including both the ID‑Sink and Context‑Cache principles.
  • w/o ID-Sink: Configuration exclude the ID‑Sink principle, retaining only the Context‑Cache principle.
  • w/o Context-Cache: Configuration exclude the ID-Sink principle, retaining only the ID-Sink principle.
The inference segment length is set to five times the training chunk length (120 frames). The performance of each configuration is evaluated on the HDTF dataset across FID, FVD, and SyncNet metrics. As detailed in Tab. 2, ablating the ID-Sink mechanism has a minimal impact on Sync-C but significantly worsens FID and FVD, indicating that suboptimal ID consistency leads to semantic distortion in generated videos. In contrast, ablating the Context-Cache mechanism results in a clear deterioration in Sync-C, accompanied by declines in FID and FVD, which can be attributed to weakened temporal consistency. To further investigate these effects, we conducted a visual analysis by uniformly sampling frames at aligned timestamps and computing difference heatmaps between adjacent frames. The results are shown in Fig. 3.
As observed, the generation results of w/o ID‑Sink configuration exhibit progressively deteriorating ID consistency over time, accompanied by a degradation in fine-grained details (e.g., the eyes in the zoomed-in views). Moreover, the overall color tone gradually degrades as inference proceeds, and the error heatmap shows high‑error regions not only on the subject but also across the background, indicating suboptimal ID consistency. Conversely, the w/o Context-Cache configuration suffers from compromised motion smoothness, characterized by abrupt motion discontinuities at chunk boundaries (see zoomed-in views). In addition, the difference heatmap displays overall higher error values distributed across the entire figure, confirming inferior temporal consistency and motion smoothness. In contrast, the Full ID-Context Cache, which integrates both principles, achieves superior performance in both identity preservation and motion smoothness. The zoom‑in details are clearly superior to those of the ablated variants. Notably, in the difference heatmaps of the results of Full ID‑Context Cache, high values are concentrated solely on dynamic facial regions (i.e., eyes and mouth), while the background and overall head structure exhibit minimal error. These visualizations confirm the crucial roles of ID-Sink and Context-Cache in enhancing identity consistency and temporal smoothness, demonstrating the effectiveness of our proposed method.
Ablation Studies on ASD. To validate the effectiveness of our proposed ASD strategy and the contribution of each loss terms, we conduct a quantitative ablation study using the following configurations:
  • Full ASD: Training with the complete ASD strategy.
  • w/o Smooth: Training without Smooth loss.
  • w/o Contrastive: Training without Contrastive loss.
  • w/o ASD: Training without the ASD strategy.
Method FID ($\downarrow$) FVD ($\downarrow$) Sync-C ($\uparrow$)
Full ASD 14.597 219.87 8.335
w/o Smooth 14.646 221.525 8.264
w/o Contrastive 15.278 211.881 8.190
w/o ASD 15.950 228.815 7.970
Table 3: Ablation results of ASD on HDTF dataset.
We report the FID, FVD, and SyncNet scores of each configuration on the HDTF dataset, as shown in Tab 3. As indicated by the results, removing the Smooth loss while retaining the Contrastive loss for teacher-student distillation leads to a deterioration in the FVD metric, whereas FID and Sync-C scores show only marginal degradation. This indicates that the Smooth loss plays a crucial role in enhancing temporal coherence and preserving the overall semantic consistency of the generated videos. Conversely, removing the Contrastive loss and retaining only the Smooth loss results in noticeable deterioration in both Sync-C and FID metrics, with minimal impact on FVD. This demonstrates that the Contrastive loss primarily contributes to improving the overall visual quality and audio-visual alignment of the student model during streaming generation. Furthermore, the complete removal of the ASD strategy shows a substantial performance drop across all three metrics compared to the Full ASD configuration. These results validate the significant efficacy of our non-streaming to streaming distillation strategy in improving both generation quality and temporal consistency.
4.4 Case Study
To provide a more intuitive comparison of the generation quality among SOTA THG methods, we present a visual analysis of a representative test case. Conditioned on the identical reference image and driving audio, we utilize each model to generate talking head videos of equal length of 120 frames. Key frames at identical timestamps are sampled for visual comparison, as illustrated in Fig. 4, with the inference runtime of backbone network for each model provided.
Figure 4: Case study of SOTA THG methods.
Visual analysis of the representative case reveals distinct performance characteristics among THG methods. Motion-space diffusion methods such as AniTalker and Ditto achieve low runtime latency, yet exhibit poor lip-sync, with visible mismatches between generated and ground truth frames at corresponding timestamps (indicated by red arrows), underscoring the limitations of their two-stage training pipelines. End-to-end models, including Sonic and EchoMimic, show improved lip alignment, though occasional inaccuracies persist. Meanwhile, Hallo suffers from noticeable artifacts in later frames, attributable to its inferior generative quality. In contrast, our model operates with minimal inference time while maintaining accurate lip synchronization under real-time streaming constraints. The generated lip motions closely align with the ground truth, and the outputs exhibit natural head movements and facial expressions. Together, these results demonstrate that our approach achieves an optimal balance between efficiency and perceptual quality, confirming the effectiveness of the proposed method.
5 Conclusion
This study addresses the critical and fundamental challenges of slow inference speed and the lack of streaming capability in existing diffusion-based talking head generation (THG) models by proposing REST, the first diffusion-based, real-time end-to-end streaming framework for audio-driven talking head generation. A key innovation of REST is the introduction of the ID-Context Cache mechanism, which efficiently caches ID-related information and historical context associated with motion continuity to enable autoregressive contextual streaming inference. Furthermore, to tackle the quality and consistency bottlenecks inherent in streaming scenarios, we further propose the asynchronous streaming distillation (ASD) strategy. This approach first employs an asynchronous noise scheduler to simulate streaming conditions, aligning the non-streaming teacher model with the streaming student model. Through the combination of teacher-student contrastive loss and motion smoothness loss, the global inter-frame dependency knowledge can be distilled from the teacher model into the student model, thereby enhancing the long-term generation quality and temporal consistency of the student model under streaming constraints. In general, REST effectively bridges autoregressive and diffusion models, achieving SOTA performance in both speed and quality under real-time streaming conditions. Extensive ablation studies further validate the effectiveness of our proposed approaches.